OpenStudy (anonymous):

2, 4, _, 10, 12, _ Will give Medal!

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

2, 4, _, 10, 12, _ Will give Medal!

OpenStudy (anonymous):

2+2=4 4+4=8 8+2+10 10+2+12 12+4=16 im not sure if that is right but it is pretty logical and a pattern